# Physics of Strong Magnetism with eXTP

### v1

@@ 3 Accretion under strong magnetic fields
Although magnetars represent one of the most extreme manifestations of strong magnetic fields in NSs, accreting X-ray pulsars (XRPs) offer a crucial avenue for exploring the physics of interaction of plasma with highly magnetized NSs. Unlike magnetars, XRPs are fueled by accretion, a fundamental astrophysical process in which matter accumulates onto a central compact object. In XRPs, where the NS has a strong magnetic field on the order of ,G, the accretion process is fundamentally shaped by magnetic forces, leading to a wide variety of observational phenomena [for reviews, see 135, 229]. The interaction of accreting plasma with a strong magnetic field gives rise to a highly complex geometrical structure in these systems (see Fig. 10). At the magnetospheric radius, where the ram pressure of the accretion disk matches the magnetic pressure, infalling matter is captured by the field lines and channeled toward the magnetic poles, where its energy is converted into X-ray emission. The structure of the emission region is determined by the mass accretion rate, manifesting itself either as a localized hot spot or as a vertically extended accretion column for luminosities below and above erg s-1, respectively [9, 136, 12]. The strong dependence of the scattering cross section on the magnetic field gives rise to highly anisotropic radiation patterns (commonly known as ‘beam patterns’), leading to intricate pulse profiles that reflect the complex interplay between the geometric and physical properties of XRPs. Unfortunately, this degeneracy has so far prevented the development of a reliable method for reconstructing the geometric configuration of such systems solely from their pulse profiles. However, the strong anisotropy in the scattering cross section induced by the magnetic field was expected to produce a high degree of polarization in the emission of XRPs [131], which, in turn, is sensitive to the geometry of the emission region.
The recent launch of the Imaging X-ray polarimetry Explorer [IXPE; 228], which was expected to resolve the aforementioned degeneracy, has instead revealed significant discrepancies between theoretical predictions and the observed polarization properties of XRPs [see review by 163]. This provides strong motivation for the next-generation X-ray observatory, eXTP, which offers a significantly larger effective area for both timing and polarimetric studies.
@@ 3.1.1 Rotating vector model and NS orientation
By now, a sample of about a dozen XRPs has been observed with IXPE: Her X-1 [43, 59, 244, 72], Cen X-3 [212], GRO J100857 [213], 4U 162667 [127], X Persei [139], Vela X-1 [53, 55], EXO 2030+375 [125], GX 3012 [188], RX J0440.9+4431/LS V +44 17 [45, 245], Swift J0243.6+6124 [120, 164], SMC X-1 [54], and 4U 153852 [117]. Although the PD may exhibit a complex evolution pattern over the pulse phase, the PA variations in most XRPs follow a much simpler trend, enabling the application of the RVM. This has enabled independent determination of the geometric configuration of at least the brightest XRPs, including the discovery of nearly aligned and nearly orthogonal rotators [163].
@@ 3.2.1 Nature of emission at very low luminosities
One of the long-standing challenges in the study of XRPs is understanding the nature of their emission in the quiescent state. This question is closely related to the verification of the proposed centrifugal inhibition of accretion, which is expected to occur at very low mass accretion rates onto highly magnetized NSs [the so-called ‘propeller effect’; 87]. The propeller effect relies on the dipole magnetic field, which decays more slowly compared to the multipole field on a large scale. As a result, it provides a straightforward approach to confirm multipole fields by comparing the magnetic field inferred by the propeller effect and other measurements, e.g., CRSFs generated close to the NS surface [105]. Depending on whether this effect occurs in nature, different sources of quiescent luminosity have been considered in the literature, with thermal cooling of the NS and continued low-level accretion being the most probable explanations [see, e.g., 209]. The key observable that can help distinguish between these two scenarios is the shape of the energy spectrum in this state: a soft, nearly thermal spectrum is expected in the absence of accretion, while a much harder spectrum is expected if accretion continues.
@@ 3.2.2 Physics of the NS atmosphere and low PD in XRPs
There are no exact self-consistent models of such atmospheres [see, however, 186, 179, 138]. However, it is possible to estimate the radiation properties of such an atmosphere using a toy model in the first approximation. In such a model, the magnetized pure hydrogen atmosphere can be represented as two isothermal plasma layers, an upper hot (10–100 keV) with an exponential temperature drop to a lower cold ( keV) semi-infinite layer [see details in 55]. Assuming that these layers are in hydrostatic equilibrium in the gravitational field of a typical NS, we can calculate the number densities of protons and electrons, the opacity coefficients, and consider the radiative transport in two polarization modes in the resulting model atmosphere.
@@
The depth of the atmosphere at which the vacuum resonance occurs for a given magnetic field depends not only on the photon energy but also on the plasma density [see details in 250]. For low-energy photons, the vacuum resonance occurs at low plasma density in the upper optically thin layers of the atmosphere. In this case, a simple exchange of fluxes between the O- and X-modes takes place, and the PA of the emergent radiation rotates by 90∘. Vacuum resonance for high-energy photons occurs in dense optically thick layers of the atmosphere and does not affect the polarization of the emergent radiation.
@@ 3.3 Critical luminosity and geometry of emission region
In observations, transitions through the critical luminosity have been suggested for several transient XRPs based on changes in their pulse profiles and the hardness evolution with luminosity. This provides an indirect method to estimate the magnetic field of a NS [e.g., 224, 156, 167]. Around the critical luminosity phase-resolved polarization properties of XRPs are expected to change dramatically [131]. These changes can be precisely measured with eXTP/PFA during observations of a transient XRP. The accurate measurement of the critical luminosity would therefore provide a crucial test for the entire paradigm of accretion onto highly magnetized NSs.
@@ 3.4 Extreme accretion: PULXs
of a 10 M⊙ black hole [BH; see reviews by 94, 49, 102, 151]. We know now of about 2000 ULXs [107, 221, 207]. Initially thought to be powered by intermediate-mass BHs ( M⊙) accreting at sub-Eddington rates [see e.g. 28], since the early 2000s a few pieces of evidence pointed towards super-Eddington accretion in X-ray binaries (XRBs) [100, 161, 238]. First of all, they are characterized by peculiar spectral states, different from those of Galactic XRBs [61, 189]. Secondly, the energy spectra of a few ULXs have shown the presence of strong outflow winds, a key prediction of the super-Eddington scenario [152, 106]. The most striking evidence in support of this scenario, however, came in 2014, when coherent pulsations were detected in the X-ray flux of M82 X-2, clearly stating that this source is an accreting NS shining at luminosities , the first pulsating ULX [PULX 6].
Since then, 5 more extragalactic PULXs emitting at luminosities far exceeding Eddington limit for a NS have been identified [6, 57, 91, 90, 26, 175, 171], and a few more accreting NSs have reached similar luminosities at least once [see 102, and references therein]. Their spin periods are characterized by the presence of huge spin-up rates () that support the super-Eddington nature of these sources [see e.g. 90, for the case of NGC 5907 ULX-1]. The super-Eddington nature is also supported by the magnitude of the orbital decay of M82 X-2 [7]. It is interesting to note that M82 X-2 was found to show a long-term spin-down trend [115]. PULXs are characterized by hard spectra [153, 66], practically indistinguishable from the X-ray spectra of many ULXs [see e.g. 220], suggesting a much larger population. Finally, at least 2 PULXs have shown mHz QPOs with similar properties in the past years [50, 88].
How PULXs can reach such high luminosities is still an open question. A (partial) solution could lie in the magnetic field of the accreting NSs powering these systems. Particularly intense dipolar magnetic fields ( G) could help overcome the Eddington limit [see, e.g., 35, 137, 203], but at the observed spin periods ( s) PULXs should constantly be in the propeller regime [87]. In the case of NGC 5907 ULX-1, it has been argued that a less intense dipolar component ( G) accompanied by a strong multipolar component near the NS surface ( G) could explain its extreme luminosities erg s [90]. A pronounced geometrical beaming of the X-ray radiation has also been invoked to explain the PULX luminosities [98, 101, 99], but although some degree of beaming must be present, an unrealistically high value of would be required to explain NGC 5907 ULX-1 luminosities [see again 90].
@@ Acknowledgements
This work is supported by the National Key R&D Program of China (2021YFA0718500). We acknowledge the support by the National Natural Science Foundation of China (Grant Nos. 12373051, and 12273028, 12173103 and 12261141691, and 12003009). RTav and RTur acknowledge the support by the Italian MUR (grant PRIN 2022 - 2022LWPEXW grant, “An X-ray view of compact objects in polarized light”, CUP C53D23001180006). VS thanks the German Research Foundation (DFG) grant WE 1312/59-1 for financial support. AAM thanks UKRI Stephen Hawking fellowship. RK aknowledges the support by UKRI STFC under the grant ST/W507891/1. S-NZ is supported by the National Natural Science Foundation of China (No. 12333007), the International Partnership Program of Chinese Academy of Sciences (No.113111KYSB20190020) and the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDA15020100). FX is supported by National Natural Science Foundation of China (grant No. 12373041 and No. 12422306), and Bagui Scholars Program (XF).

### v2

@@ 3 Accretion under strong magnetic fields
Although magnetars represent one of the most extreme manifestations of strong magnetic fields in NSs, accreting X-ray pulsars (XRPs) offer a crucial avenue for exploring the physics of interaction of plasma with highly magnetized NSs. Unlike magnetars, XRPs are fueled by accretion, a fundamental astrophysical process in which matter accumulates onto a central compact object. In XRPs, where the NS has a strong magnetic field on the order of ,G, the accretion process is fundamentally shaped by magnetic forces, leading to a wide variety of observational phenomena [for reviews, see 45, 193]. The interaction of accreting plasma with a strong magnetic field gives rise to a highly complex geometrical structure in these systems (see Fig. 12). At the magnetospheric radius, where the ram pressure of the accretion disk matches the magnetic pressure, infalling matter is captured by the field lines and channeled toward the magnetic poles, where its energy is converted into X-ray emission. The structure of the emission region is determined by the mass accretion rate, manifesting itself either as a localized hot spot or as a vertically extended accretion column for luminosities below and above erg s-1, respectively [194, 195, 196]. The strong dependence of the scattering cross section on the magnetic field gives rise to highly anisotropic radiation patterns (commonly known as ‘beam patterns’), leading to intricate pulse profiles that reflect the complex interplay between the geometric and physical properties of XRPs. Unfortunately, this degeneracy has so far prevented the development of a reliable method for reconstructing the geometric configuration of such systems solely from their pulse profiles. However, the strong anisotropy in the scattering cross section induced by the magnetic field was expected to produce a high degree of polarization in the emission of XRPs [197], which, in turn, is sensitive to the geometry of the emission region.
The recent launch of the Imaging X-ray polarimetry Explorer [IXPE; 198], which was expected to resolve the aforementioned degeneracy, has instead revealed significant discrepancies between theoretical predictions and the observed polarization properties of XRPs [see review by 52]. This provides strong motivation for the next-generation X-ray observatory, eXTP, which offers a significantly larger effective area for both timing and polarimetric studies.
@@ 3.1.1 Rotating vector model and NS orientation
| Source Name | Period (s) | Orbital Period (d) | (keV) | (deg) | (deg) | IXPE Reference |
| Her X-1 | 1.24 | 1.7 | 37 | , 90 30 | , | [200, 201, 202, 98] |
| Cen X-3 | 4.8 | 2.09 | 28 | 70.2 (fixed) | 16.4 1.3 | [192] |
| GRO J100857 | 93.5 | 249.5 | 78 | 130 3 | 74 2 | [203] |
| 4U 162667 | 7.7 | 0.02875 | 37, 61? | - | - | [204] |
| X Persei | 837.67 | 250.3 | 29 | 162 12 | 90 15 | [205] |
| Vela X-1 | 283 | 8.96 | 25, 53 | - | - | [206, 207] |
| EXO 2030+375 | 41.31 | 46.02 | 36/63? | [208] | ||
| GX 3012 | 696.0 | 41.59 | 50 | 135 17 | 43 12 | [209] |
| RX J0440.9+4431 | 202.5 | 155.0 | 32? | 108 2 | 48 1 | [210, 211] |
| Swift J0243.6+6124 | 9.87 | 28.3 | 146 | [212, 213] | ||
| SMC X-1 | 0.717 | 3.892 | - | [214] | ||
| 4U 153852 | 526.42 | 3.73 | 22 | [215] | ||
| 4U 1907+09 | 438.0 | 8.37 | 18, 36 | - | - | [216] |
By now, a sample of about a dozen XRPs has been observed with IXPE (see Table 1). Although the PD may exhibit a complex evolution pattern over the pulse phase, the PA variations in most XRPs follow a much simpler trend, enabling the application of the RVM. This has enabled independent determination of the geometric configuration of at least the brightest XRPs, including the discovery of nearly aligned and nearly orthogonal rotators [52].
@@ 3.2.1 Nature of emission at very low luminosities
One of the long-standing challenges in the study of XRPs is understanding the nature of their emission in the quiescent state. This question is closely related to the verification of the proposed centrifugal inhibition of accretion, which is expected to occur at very low mass accretion rates onto highly magnetized NSs [the so-called ‘propeller effect’; 220]. The propeller effect relies on the dipole magnetic field, which decays more slowly compared to the multipole field on a large scale. As a result, it provides a straightforward approach to confirm multipole fields by comparing the magnetic field inferred by the propeller effect and other measurements, e.g., CRSFs generated close to the NS surface [221]. Depending on whether this effect occurs in nature, different sources of quiescent luminosity have been considered in the literature, with thermal cooling of the NS and continued low-level accretion being the most probable explanations [see, e.g., 222]. The key observable that can help distinguish between these two scenarios is the shape of the energy spectrum in this state: a soft, nearly thermal spectrum is expected in the absence of accretion, while a much harder spectrum is expected if accretion continues.
@@ 3.2.2 Physics of the NS atmosphere and low PD in XRPs
There are no exact self-consistent models of such atmospheres [see, however, 225, 50, 49]. However, it is possible to estimate the radiation properties of such an atmosphere using a toy model in the first approximation. In such a model, the magnetized pure hydrogen atmosphere can be represented as two isothermal plasma layers, an upper hot (10–100 keV) with an exponential temperature drop to a lower cold ( keV) semi-infinite layer [see details in 207]. Assuming that these layers are in hydrostatic equilibrium in the gravitational field of a typical NS, we can calculate the number densities of protons and electrons, the opacity coefficients, and consider the radiative transport in two polarization modes in the resulting model atmosphere.
@@
The depth of the atmosphere at which the vacuum resonance occurs for a given magnetic field depends not only on the photon energy but also on the plasma density [see details in 68]. For low-energy photons, the vacuum resonance occurs at low plasma density in the upper optically thin layers of the atmosphere. In this case, a simple exchange of fluxes between the O- and X-modes takes place, and the PA of the emergent radiation rotates by 90∘. Vacuum resonance for high-energy photons occurs in dense optically thick layers of the atmosphere and does not affect the polarization of the emergent radiation.
@@ 3.3 Critical luminosity and geometry of emission region
In observations, transitions through the critical luminosity have been suggested for several transient XRPs based on changes in their pulse profiles and the hardness evolution with luminosity. This provides an indirect method to estimate the magnetic field of a NS [e.g., 228, 229, 230]. Around the critical luminosity phase-resolved polarization properties of XRPs are expected to change dramatically [197]. These changes can be precisely measured with eXTP/PFA during observations of a transient XRP. The accurate measurement of the critical luminosity would therefore provide a crucial test for the entire paradigm of accretion onto highly magnetized NSs.
@@ 3.4 Extreme accretion: PULXs
of a 10 M⊙ black hole [BH; see reviews by 233, 234, 235, 236]. We know now of about 2000 ULXs [237, 238, 239]. Initially thought to be powered by intermediate-mass BHs ( M⊙) accreting at sub-Eddington rates [see e.g. 240], since the early 2000s a few pieces of evidence pointed towards super-Eddington accretion in X-ray binaries (XRBs) [241, 242, 243]. First of all, they are characterized by peculiar spectral states, different from those of Galactic XRBs [244, 245]. Secondly, the energy spectra of a few ULXs have shown the presence of strong outflow winds, a key prediction of the super-Eddington scenario [246, 247]. The most striking evidence in support of this scenario, however, came in 2014, when coherent pulsations were detected in the X-ray flux of M82 X-2, clearly stating that this source is an accreting NS shining at luminosities , the first pulsating ULX [PULX 248].
Since then, 5 more extragalactic PULXs emitting at luminosities far exceeding Eddington limit for a NS have been identified [248, 249, 250, 251, 252, 253, 254], and a few more accreting NSs have reached similar luminosities at least once [see 235, and references therein]. Their spin periods are characterized by the presence of huge spin-up rates () that support the super-Eddington nature of these sources [see e.g. 251, for the case of NGC 5907 ULX-1]. The super-Eddington nature is also supported by the magnitude of the orbital decay of M82 X-2 [255]. It is interesting to note that M82 X-2 was found to show a long-term spin-down trend [256]. PULXs are characterized by hard spectra [257, 258], practically indistinguishable from the X-ray spectra of many ULXs [see e.g. 259], suggesting a much larger population. Finally, at least 2 PULXs have shown mHz QPOs with similar properties in the past years [260, 261].
How PULXs can reach such high luminosities is still an open question. A (partial) solution could lie in the magnetic field of the accreting NSs powering these systems. Particularly intense dipolar magnetic fields ( G) could help overcome the Eddington limit [see, e.g., 262, 263, 264], but at the observed spin periods ( s) PULXs should constantly be in the propeller regime [220]. In the case of NGC 5907 ULX-1, it has been argued that a less intense dipolar component ( G) accompanied by a strong multipolar component near the NS surface ( G) could explain its extreme luminosities erg s [251]. A pronounced geometrical beaming of the X-ray radiation has also been invoked to explain the PULX luminosities [265, 266, 267], but although some degree of beaming must be present, an unrealistically high value of would be required to explain NGC 5907 ULX-1 luminosities [see again 251].
@@ Acknowledgements
This work is supported by China’s Space Origins Exploration Program and the National Key R&D Program of China (2021YFA0718500). We acknowledge the support by the National Natural Science Foundation of China (Grant Nos. 12373051, and 12273028, 12173103 and 12261141691, and 12003009). RTav and RTur acknowledge the support by the Italian MUR (grant PRIN 2022 - 2022LWPEXW grant, “An X-ray view of compact objects in polarized light”, CUP C53D23001180006). VS thanks the German Research Foundation (DFG) grant WE 1312/59-1 for financial support. AAM thanks UKRI Stephen Hawking fellowship. RK aknowledges the support by UKRI STFC under the grant ST/W507891/1. S-NZ is supported by the National Natural Science Foundation of China (No. 12333007), the International Partnership Program of Chinese Academy of Sciences (No.113111KYSB20190020) and the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDA15020100). FX is supported by National Natural Science Foundation of China (grant No. 12373041 and No. 12422306), and Bagui Scholars Program (XF). NB is supported by STFC (grant code: ST/Y001060/1).
Interest Conflict: The authors declare that they have no conflict of interest.